John has twice as much money invested at 5% simple annual interest as she does at 4%. If her yearly income from the two investments is $245, how much does she have at each rate?
at 4% she has ___?
at 5% she has ___?

Answer by mananth(16946)

At 4 % ---------------x
at 5% ----------------2x
4%x+5%*2x=245
multiply by 100
4x+10x=24500
14x=24500
/14
x=1750
$1750 at 4%
$3500 at 5%